OpenStudy (anonymous):

A bakery made 28 cherry pies, using 111 cherries for each pie. They threw away 31 cherries that were bad. If they used all the cherries they had, how many cherries did they start off with? A. 170 B. 3,139 C. 3,108 D. 3,077

OpenStudy (texaschic101):

28 cherry pies, 111 cherries per pie = 28 * 111 = 3108 cherries used 31 were bad....3108 - 31 = 3077

OpenStudy (anonymous):

Number of pies (28) x Number of cherries in each pie (111) = 3108 cherries used for pies 3108 + 31(bad cherries) = 3139 (b)